# Forecasting Galaxy Cluster H i Mass Recovery with CHIME at Redshifts z=1 and 2 via the IllustrisTNG Simulations

### v1

@@ Acknowledgements
The authors thank Frank van den Bosch and Héctor Arce for a fruitful discussion that helped this paper take its initial shape. AP thanks Seth Siegel, Simon Foreman, and Richard Shaw, along with the rest of the Canadian Hydrogen Intensity Mapping Experiment Collaboration, for help generating CHIME-like beams, assessing simulation suitability, and useful comments that improved the manuscript.
@@
Data used in this paper are publicly available as part of TNG300 simulation (Pillepich et al., 2018; Springel et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) from the IllustrisTNG project [^5] https://www.tng-project.org/data/downloads/TNG300-1/ (Nelson et al., 2019). Additionally, our beams, catalogs, tiled intensity maps, and analysis code are available on GitHub [^6] https://github.com/avapolzin/CHIMExIllustrisTNG – the repository will be made public on paper acceptance..
@@ Appendix A Making a tiled map from the simulation volume
We iterate through different projections of the simulation volume in order to construct a larger map that better replicates the angular coverage of real CHIME data.
@@ B.1 Mass-selected galaxy clusters
In generating the mass-selected galaxy cluster sample, we start by selecting those haloes with $M_{200c}\geq 10^{14}\,M_{\odot}$ within the $z=0$ group catalog and then track their central subhalo back from $z=0$ to $z=2$ or $1$, using the Main Progenitor Branch of the TNG300 SubLink merger trees via the API. For these (sub)haloes, we store $M_{200c}$ at $z=0$, as well as the group ID and position of the massive progenitor halo at $z=1,2$.
@@
Before any permutations of the simulation volume, the mass-selected galaxy cluster sample contains 276 haloes at $z=2$ and 277 haloes at $z=1$, down from 280 haloes with $M_{200c}\geq 10^{14}\,M_{\odot}$ at $z=0$.
@@ B.2 Radius-selected galaxy clusters
Selecting clusters by their radius at $z=1$ or $2$ is somewhat more straightforward. We define radius-selected galaxy clusters at each redshift as haloes with $r_{200c}\geq 0.75\,h^{-1}$ Mpc. (See also D’Aloisio et al. 2009 for a more in-depth discussion of proto-cluster and cluster structure.)

### v2

@@ Acknowledgements
The authors thank Frank van den Bosch and Héctor Arce for a fruitful discussion that helped this paper take its initial shape and thank the anonymous referee for comments and suggestions that improved the manuscript. AP thanks Seth Siegel, Simon Foreman, and Richard Shaw, along with the rest of the Canadian Hydrogen Intensity Mapping Experiment Collaboration, for help generating CHIME-like beams, assessing simulation suitability, and useful comments that enhanced this work.
@@
Data used in this paper are publicly available as part of TNG300 simulation (Pillepich et al., 2018; Springel et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) from the IllustrisTNG project [^5] https://www.tng-project.org/data/downloads/TNG300-1/ (Nelson et al., 2019). Additionally, our beams, catalogs, tiled intensity maps, and analysis code are available on GitHub [^6] https://github.com/avapolzin/CHIMExIllustrisTNG.
@@ Appendix A Making a tiled map from the simulation volume
The volume is projected along a single axis, with the galaxy’s luminosity distance defined by $d_{L}(z)$ plus its location along the collapsed, projected axis of the volume. We iterate through different projections of the simulation volume in order to construct a larger map that better replicates the angular coverage of real CHIME data.
@@ B.1 Mass-selected galaxy clusters
In generating the mass-selected galaxy cluster sample, we start by selecting those haloes with $M_{200}\geq 10^{14}\,M_{\odot}$ within the $z=0$ group catalog and then track their central subhalo back from $z=0$ to $z=2$ or $1$, using the Main Progenitor Branch of the TNG300 SubLink merger trees via the API. For these (sub)haloes, we store $M_{200}$ at $z=0$, as well as the group ID and position of the massive progenitor halo at $z=1,2$.
@@
Before any permutations of the simulation volume, the mass-selected galaxy cluster sample contains 276 haloes at $z=2$ and 277 haloes at $z=1$, down from 280 haloes with $M_{200}\geq 10^{14}\,M_{\odot}$ at $z=0$.
@@ B.2 Radius-selected galaxy clusters
Selecting clusters by their radius at $z=1$ or $2$ is somewhat more straightforward. We define radius-selected galaxy clusters at each redshift as haloes with $R_{200}\geq 0.75\,h^{-1}$ cMpc. (See also D’Aloisio et al. 2009 for a more in-depth discussion of proto-cluster and cluster structure.)